OpenStudy (anonymous):

Math Help? Simplify: 5m^2 n^3·mn^2(4m - n)

OpenStudy (anonymous):

Simplify 5m2n3 - mn2•(4m-n)

OpenStudy (anonymous):

5m2n3 - 4m2n2 + mn3 = mn2 • (5mn - 4m + n)

OpenStudy (anonymous):

mn2 • (5mn - 4m + n) = 0

OpenStudy (anonymous):

mn2•(5mn-4m+n) = 0

OpenStudy (anonymous):

What are we looking for?

OpenStudy (anonymous):

Or just simplifying the expression as a whole

OpenStudy (anonymous):